OpenStudy (anonymous):

Which of the following measurements is equal to 550000 µL? 5.5L 55 cL 5500 mL 5500000 nL

OpenStudy (anonymous):

55 cL

OpenStudy (anonymous):

1 L = 100 cL 0.55 L = x cL 0.55 * 100 = 55 cL 55 cL is true.... Hope i Helped....

OpenStudy (anonymous):

1 uL = 1.0*10^-6 L so 10 uL = 1.0*10^-5 L therefore 550000 = 5.5*10^-1 L and 5.5*10^-1 L = 0.55 L 0.55 L = 550 mL = 55 centiliters